You have five(5) 10.0 Farad capacitors. Find the total capacitance if the capacitors are connected in series.

1
Expert's answer
2022-03-21T12:42:04-0400

The net capacitance

"\frac{1}{C_{net}}=\frac{1}{C_{1}}+...+\frac{1}{C_{5}}=\frac{5}{C}"

"C_{net}=\frac{10.0\:\rm F}{5}=2.00\:\rm F"